jo-ann owns jewerly store. She used 150 pairs of earrings as her sample of the different design. The population standard deviation of the price of the earrings is 65 pesos. suppose that jo-ann wants a 98% level of confidence to determine the mean of price all her earrings she is selling. Compute for length of the confidence retrival


1
Expert's answer
2022-03-13T14:56:09-0400

The confidence interval for the mean is

"\left( \bar{x}-z_{1-\frac{\alpha}{2}}\frac{\sigma}{\sqrt{n}},\bar{x}+z_{1-\frac{\alpha}{2}}\frac{\sigma}{\sqrt{n}} \right)"

from which the length of the interval is

"2z_{1-\frac{\alpha}{2}}\frac{\sigma}{\sqrt{n}}=2\cdot 2.3263\cdot \frac{65}{\sqrt{150}}=24.6924"